OpenStudy (anonymous):

17. Given a slope of 13 and a y-intercept of -5, write the equation of the line in slope-intercept form. A. y = 13 - 5x B. y = 13x - 5 C. y = 13x + 5

OpenStudy (mokeira):

that means x=0 y=mx+c \[m=\frac{ \Delta y }{ \Delta x }\] \[13=\frac{ y--5 }{x-0 }\] cross multiply \[13(x)=y+5\] arrange it accordingly \[y=13x-5\]

OpenStudy (anonymous):

:( @Mokeira good job